Find z that has an area
(a) 0.9505 to its left
(b) 0.05 to its left
(c) 0.6915 to its right
(d) 0.1587 to its right
1
Expert's answer
2019-08-29T07:34:35-0400

https://www.sjsu.edu/faculty/gerstman/EpiInfo/z-table.htmIt is table of standard normal distribution (more precisely, in this table "Z(z)=\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^z e^{-\frac{x^2}{2}}dx" )So having this table we can write:a)"Z(z)=0.9505" if "z=1.65" b)Since graph of standard normal distribution is symmetric with respect to line "x=0" , for every "u\ge 0" we have "\frac{1}{\sqrt{2\pi}}\int\limits_{-u}^0 e^{-\frac{x^2}{2}}dx=\frac{1}{\sqrt{2\pi}}\int\limits_0^u e^{-\frac{x^2}{2}}dx". (1)Let "z" such number that "Z(z)=0,05" . So "0.5=\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^0 e^{-\frac{x^2}{2}}dx=\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^z e^{-\frac{x^2}{2}}dx+\frac{1}{\sqrt{2\pi}}\int\limits_z^0 e^{-\frac{x^2}{2}}dx=" "=0.05+\frac{1}{\sqrt{2\pi}}\int\limits_z^0 e^{-\frac{x^2}{2}}dx" . That is "\frac{1}{\sqrt{2\pi}}\int\limits_z^0 e^{-\frac{x^2}{2}}dx=0.45=\frac{1}{\sqrt{2\pi}}\int\limits_0^{-z} e^{-\frac{x^2}{2}}dx" by property (1). We have "Z(-z)=\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^{-z} e^{-\frac{x^2}{2}}dx=\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^0 e^{-\frac{x^2}{2}}dx+\frac{1}{\sqrt{2\pi}}\int\limits_0^{-z} e^{-\frac{x^2}{2}}dx=0.95" . From table we can see that "-z\approx 1.64" , so "z\approx -1.64" c)We need "z" such "\frac{1}{\sqrt{2\pi}}\int\limits_z^{+\infty} e^{-\frac{x^2}{2}}dx=0.6915" . Since "1=\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^{+\infty} e^{-\frac{x^2}{2}}dx=\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^z e^{-\frac{x^2}{2}}dx+\frac{1}{\sqrt{2\pi}}\int\limits_z^{+\infty} e^{-\frac{x^2}{2}}dx=" "=Z(z)+0.6915" , so "Z(z)=1-0.6915=0.3085" Applying the same argument as in b). Since "0.3085=0.5-0.1915" , we have "Z(-z)=0.5+0.1915=0.6915" . So from the table we see that "-z=0.50" and so "z=-0.50" d)Applying the same argument as in c), we have "Z(z)=1-0.1587=0.8413" . From the table we see that "z=1.00"